gelphielvr:

(algebra) How would I solve an equation like this? 2/5x -5 = -1x +2 (I'm looking how to solve it in general not just for the answer.)

gelphielvr:

@timmyspu @yomomalikebracobama

gelphielvr:

I know how to do it with a calculator but were not allowed to use those for our test ... so if anyone knows the steps pls help

Graysonunderwoods:

damn i dont even know hot to solve that with out a mathamic calculator

gelphielvr:

@graysonunderwoods wrote:
damn i dont even know hot to solve that with out a mathamic calculator
I think these are the steps but I don't understand how the duck this happened. 1. !x + 2/5x = 2 + 5 2. 1x + 2/5x = 7 3. ? 4. 5x+2x/5 5. 7x/5 = 7 and then x is 5 but idk how that happened

Graysonunderwoods:

i believe for the most part thats right google ismightly helpful,so i just looked it up

Luigi0210:

So you're given \(\large \frac{2}{5} x - 5 = -x+2 \) What they did in your example is add x to both sides \(\large \frac{2}{5}x +x -5 = -x+x +2 \) The -x and x cancel. Then you can convert 1x into 5/5 x, which is the same as 1x \(\large \frac{2}{5} x+ \frac{5}{5} x -5 = 2 \) Combine the x's, and then add 5 to both sides \(\large \frac{7}{5} x = 7 \) Now multiply by 5 to get rid of the fraction \(\large \frac{7*5}{5} x = 7*5 \) \(\large 7x= 35 \) Now just divide by 7 \(\large \frac{7x}{7} =\frac{35}{7} \) \(\large x= 5\)